OpenStudy (decarr432):

David is rowing a boat upstream. The river is flowing at a speed of 2 miles per hour. David starts rowing at a speed of 6 miles per hour, but as he gets tired his speed decreases (at a rate of 1 mile per hour, every hour). Which equation represents the speed of the boat for x hours spent rowing?

OpenStudy (decarr432):

A.y=x+4 B.y=6-x C.y=4-x D.y=6-2x E.y=x-4

OpenStudy (decarr432):

@zepdrix

OpenStudy (thecuteone25):

@Decarr432

OpenStudy (decarr432):

I think its be because he starts at six but decrees is by 1x or just x

OpenStudy (thecuteone25):

I think your right

OpenStudy (thecuteone25):

yeah you are sorry